Procedure:

  • remember to people on the summit to:
    • switch on the main cooling
    • switch on the Neslab
    • check 110V is already plugged in
    • plug in the 208V power

  • our software:
    • set to lab-mode
    • stop AO arbitrator
    • power on and load program
    • stop Adsec Arbitrator
    • set the shell using terminal: print, fsm_set_flat(/NO_FLAT)
    • start Adsec Arbitrator
    • disable HO Offload
    • send to teminal. print, setpistoncurrent(20.0)

Test

LV = little volcano (order of 10mN)

  • act 366 (es 2011-11-18 14:50:00) (LV: to be removed from ICC)
  • act 500 (14:59:00) (LV: to be removed from ICC)

  • start slewing: ???
    • bunch of 2x2 acts: 8, 9, 71, 56 (LV: to be removed from ICC)
  • end slewing: 2011-11-18 15:19:00

  • it seems that astigmatism is oscillating (2011-11-18 15:25:10)
  • act 515 (2011-11-18 15:34:21) (LV: to be removed from ICC)

  • inserted in wo_icc list: 366 , 500 , 8, 9, 71, 56 , 515 (new elec.txt configuration: 20111118_173000)

  • start slewing: 2011-11-18 17:55:47
    • nothing relevant
  • stop slewing: 2011-11-18 17:44:56

  • astigmatism problem is still there
  • next test session on Monday
-- ArmandoRiccardi - 18 Nov 2011
Topic revision: r2 - 18 Nov 2011, ArmandoRiccardi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback